What is the molarity of a 250.0 mL aqueous solution that contains 18.7 g of AgNO3
1
Expert's answer
2020-09-21T06:55:26-0400

СM=n/V

n=m/M

M(AgNO3)=170 g/mol

n (AgNO3) = 18.7/170=0.11 mol

СM=0.11/0.25=0.44 M